OpenStudy (anonymous):

Verify the identity. 4csc2x = 2csc^2xtan x

OpenStudy (slaaibak):

4csc2x = 4/sin2x = 4/(2sinxcosx) = 4sinx/(2sin^2xcosx) = 2tanxcsc^2x

OpenStudy (anonymous):

How did you get 4/(2sinxcosx)?

OpenStudy (anonymous):

@slaaibak

OpenStudy (slaaibak):

sin2x = 2sinxcosx It's a double angle identity

OpenStudy (anonymous):

Ohh ok

OpenStudy (anonymous):

Thanks so much

OpenStudy (slaaibak):

it's a pleasure :)